OpenStudy (anonymous):

Subtract. –7 – 12 A. –19 B. –5 C. 5 D. 19

OpenStudy (alexandervonhumboldt2):

-7-12=-(7+12)=-19

OpenStudy (anonymous):

-7 - 12 = -19